Convert the interest rate J2=9% to J12 equivalent
1
Expert's answer
2020-06-24T18:12:10-0400

Jm = i*m

Nominal interest rate = 9% per annum.

J2=9%

J1 = "J_{2} \times 2 = 9\% \times 2 = 18\%"

J12= "J_{1}/12 = 18\% / 12 = 1.5\%"

J2 is an interest rate of 9% per annum implying that it is compounded semi-annually. To convert it to J12 requires the conversion to J1 to get the yearly interest rate compounded annually which in this case is equal to 18%. To get J12, the J1 interest is divided by 12 months to get 1.5% which means that and interest rate of 1.5 percent is charged on a monthly basis.